OpenStudy (anonymous):

3x^2=21 What are the solutions?

OpenStudy (anonymous):

x= +2.64575 or -2.64575

OpenStudy (anonymous):

so the answer is x=2.65, -2.65?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Divide both sides by 3, you get \(x^2=7\), that implies that x is either \(\sqrt {7}\) or \(-\sqrt{7}\).

OpenStudy (anonymous):

thank you